Question 8
ProbabilityA bag contains 5 red balls, 4 blue balls and 3 green balls. Two balls are drawn successively without replacement. What is the probability that the two balls are of different colours?
- A.7/11
- B.35/66
- C.47/66
- D.19/33
Show the answer and explanation
Answer: C (47/66)
Explanation
P(same colour) = (5×4 + 4×3 + 3×2)/(12×11) = 38/132 = 19/66, so P(different) = 1 − 19/66 = 47/66.